FSA is a privately held firm. As an analyst trying to determine the value of FSA’s common stock and bonds, you have estimated the market value of the firm’s assets to be $1 million and the standard deviation of the asset return to be .3. The debt of FSA, which consists of zero-coupon bank loans, will come due one year from now at its face value of $1 million. Assuming that the risk-free rate is 5 percent, use the Black-Scholes Model to estimate the value of the firm’s equity and debt.
Answer #1

Let's use the Black Scholes model to value equity as call option.

value of Call = S × N(d1)-K × e (-RfxT) × N(d2))

Where

d1 =

d2 = d1 - sigma sqrt{T}

S = Market Value of the firm's asset = $ 1 million

K = Strike price = Face Value of debt outstanding = $ 1 million

T = Life of the option = Due time of zero-coupon debt = 1 year

sigma= standard deviation of the asset return = 0.3

r = risk free rate = 5%

lng) + (0.05 + 0.32/2) × 1 0.3V1 di = = 0.32

d2 = 0.32 - 0.3 x sqrt{1} = 0.02

N(d1) = 0.62

N(d2) = 0.51

Value of equity as call option

(-RfxT)

1 × 0.32-1 × e(-0.05×1)

=$ 0.1423 mn

Value of debt = Value of asset - Value of equity = S - E = 1 - 0.1423 = $ 0.8577 mn